Circular No. 7488

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2000 R1 (SHOEMAKER-LEVY 5) The LINEAR team reported that one of their objects observed on Sept.

6 was cometary, and T. Spahr and the undersigned noted that this was ∼ 0^{o}.6 northwest of the prediction for P/1991 T1 = 1991z = 1991 XXII (MPC 29882, 39661), corresponding to ΔT = −1.4 day. Following a request from the Central Bureau, D. Balam reports that images taken with the 1.82-m Plaskett telescope of the National Research Council of Canada show this object to have a well-condensed coma and a fan-shaped tail extending 30′′ in p.a. 250^{o}; his R magnitude given below was obtained in a 10′′ aperture.

2000 UT α_{2000} δ_{2000} m_{1} Observer
Sept.6.41220 4^{h}52^{m}12.^{s}52 +26^{o}02′23′′.2 18.5 LINEAR
6.42324 4 52 13.70 +26 02 29.3 18.7  _{”}
6.43428 4 52 14.81 +26 02 35.7 18.6  _{”}
6.44532 4 52 15.88 +26 02 42.5 18.6  _{”}
6.45637 4 52 17.10 +26 02 49.8 18.3  _{”}
7.40247 4 53 55.55 +26 12 03.5 17.7 Balam
7.41707 4 53 57.01 +26 12 12.0  _{”}
7.46828 4 54 02.27 +26 12 42.6  _{”}

Orbital elements from 42 observations, 1991–2000 (mean residual 1′′.0): Epoch = 2000 Aug. 4.0 TT

T = 2000 Aug. 17.0846 TT  ω = 6.^{o}2331
e = 0.529360  Ω = 29.6904 }2000.0
q = 1.988623 AU  i = 11.7706

a = 4.225360 AU n^{o} = 0.1134772 P = 8.686 years

2000TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Sept. 3 4^{h}46.^{m}19 +25^{o}28.^{′}7 1.766 1.994 87.^{o}4 30.^{o}4 17.7
13 5 03.33 +27 06.4 1.678 2.003  93.1 30.1 17.6
23 5 18.75 +28 41.6 1.593 2.015  99.3 29.4 17.6
Oct. 3 5 31.93 +30 16.2 1.513 2.031 106.1 28.2 17.5
13 5 42.34 +31 51.6 1.438 2.051 113.5 26.5 17.5
23 5 49.44 +33 28.6 1.372 2.074 121.6 24.1 17.4

SUPERNOVA 1999gv IN ANONYMOUS GALAXY

Corrigendum. Gal-Yam and Maoz provide a revised position for the hostgalaxy center (IAUC 7405, line 11): α = 16^{h}36^{m}02^{s}.15, δ = +66^{o}12′34′′.1.

2000 September 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green